# AgentClinic: a multimodal agent benchmark to evaluate AI in simulated clinical environments

### v3

@@ Abstract
Diagnosing and managing a patient is a complex, sequential decision making process that requires physicians to obtain information—such as which tests to perform—and to act upon it. Recent advances in artificial intelligence (AI) and large language models (LLMs) promise to profoundly impact clinical care. However, current evaluation schemes overrely on static medical question-answering benchmarks, falling short on interactive decision-making that is required in real-life clinical work. Here, we present AgentClinic: a multimodal benchmark to evaluate LLMs in their ability to operate as agents in simulated clinical environments. In our benchmark, the doctor agent must uncover the patient’s diagnosis through dialogue and active data collection. We present two open medical agent benchmarks: a multimodal image and dialogue environment, AgentClinic-NEJM, and a dialogue-only environment, AgentClinic-MedQA. We embed cognitive and implicit biases both in patient and doctor agents to emulate realistic interactions between biased agents. We find that introducing bias leads to large reductions in diagnostic accuracy of the doctor agents, as well as reduced compliance, confidence, and follow-up consultation willingness in patient agents. Evaluating a suite of state-of-the-art LLMs, we find that several models that excel in benchmarks like MedQA are performing poorly in AgentClinic-MedQA. We find that the LLM used in the patient agent is an important factor for performance in the AgentClinic benchmark. We show that both having limited interactions as well as too many interaction reduces diagnostic accuracy in doctor agents. The code and data for this work is publicly available at AgentClinic.github.io.
Figure 1: Composing and running language agents in AgentClinic. (Left) Agents are composed of several elements in AgentClinic: an LLM backbone, context, a role, and potential biases. Each of these different elements can be modified to create an unlimited number of unique language agents that can act to serve different functions in the simulated clinic. (Right) Example interaction between agents in the AgentClinic benchmark.
Introduction
@@
One of the primary goals in Artificial Intelligence (AI) is to build interactive systems that are able to solve a wide variety of problems. The field of medical AI inherits this aim, with the hope of making AI systems that are able to solve problems which can improve patient outcomes. Recently, many general-purpose large language models (LLMs) have demonstrated the ability to solve hard problems, some of which are considered challenging even for humans[1]. Among these, LLMs have quickly surpassed the average human score on the United States Medical Licensing Exam (USMLE) in a short amount of time, from 38.1% in September 2021[2] to 90.2% in November 2023[3] (human passing score is 60%, human expert score is 87%[4]). While these LLMs are not designed nor designed to replace medical practitioners, they could be beneficial for improving healthcare accessibility and scale for the over 40% of the global population facing limited healthcare access[5] and an increasingly strained global healthcare system[6].
@@
However, there still remain limitations to these systems that prevent their application in real-world clinical environments. Recently, LLMs have shown the ability to encode clinical knowledge[7, 8], retrieve relevant medical texts[9, 10], and perform accurate single-turn medical question-answering[11, 3, 12, 13]. However, clinical work is a multiplexed task that involves sequential decision making, requiring the doctor to handle uncertainty with limited information and finite resources while compassionately taking care of patients and obtaining relevant information from them. This capability is not currently reflected in the static multiple choice evaluations (that dominate the recent literature) where all the necessary information is presented in a case vignettes and where the LLM is tasked to answer a question, or to just select the most plausible answer choice for a given question.
@@
In this work, we introduce AgentClinic, an open-source multimodal agent benchmark for simulating clinical environments. We improve upon prior work by simulating many parts of the clinical environment using language agents in addition to patient and doctor agents. Through the interaction with a measurement agent, doctor agents can perform simulated medical exams (e.g. temperature, blood pressure, EKG) and order medical image readings (e.g. MRI, X-ray) through dialogue. We also support the ability for agents to exhibit 24 different biases that are known to be present in clinical environments. Furthermore, our evaluation metrics go beyond diagnostic accuracy by giving emphasis to the patient agents with measures like patient compliance and consultation ratings. Our key contributions are summarized as follows:
@@
    An open-source clinical agent benchmark with automated feedback mechanisms, including 107 patient agents with unique family histories, lifestyle habits, age categories, and diseases. This also includes an agent-based system for providing simulated medical exams (e.g. temperature, blood pressure, EKG) based on realistic disease test findings. We also present 15 multimodal agents which require an understanding of both image and text.
@@
    Results of the diagnostic accuracy of six language models on AgentClinic-MedQA: GPT-4, GPT-4o, Mixtral-8x7B, GPT-3.5, Llama 3 70B-instruct, and Llama 2 70B-chat. We also evaluate three language models on the multimodal AgentClinic-NEJM benchmark: GPT-4o, GPT-4-turbo, and GPT-4-vision-preview.
@@
    A system for incorporating complex biases that can affect the dialogue and decisions of patient and doctor agents. We present results on diagnostic accuracy and patient perception for agents that are affected by cognitive and implicit biases with Mixtral-8x7B and GPT-4. We find that doctor and patient biases can lower diagnostic accuracy, affect the patient’s willingness to follow through with treatment (compliance), reduce patient’s confidence in their doctor, and lower willingness for follow-up consultations.
@@
    We find that the language model powering the patient agent is critical for diagnostic success in this benchmark. We also show that doctor agents excel in a specific range of conversation turns, while more or less interactions reduces their diagnostic accuracy.
  5. 5.
    A clinical reader study to annotate how realistic the simulated patient and doctor interactions are, as well as how well the measurement agent represents the medical tests.
Figure 2: Process of conversion from USMLE question, to AgentClinic-MedQA Objective Structured Clinical Examination (OSCE) template, to building a patient agent that is powered by a large language model (LLM).
AgentClinic: a multimodal agent benchmark for simulating clinical environments
In this section we describe AgentClinic, which uses language agents to simulate the clinical environment.
@@
Four language agents are used in the AgentClinic benchmark: a patient agent, doctor agent, measurement agent, and a moderator. Each language agent has specific instructions and is provided unique information that is only available to that particular agent. These instructions are provided to an LLM which carries out their particular role. The doctor agent serves as the model whose performance is being evaluated, and the other three agents serve to provide this evaluation. The language agents are described in detail below.
Patient agent
Measurement agent
Doctor agent
The doctor agent serves as the primary object that is being evaluated. This agent is initially provided with minimal context about what is known about the patient as well as a brief objective (e.g. "Evaluate the patient presenting with chest pain, palpitations, and shortness of breath"). They are then instructed to investigate the patients symptoms via dialogue and data collection to arrive at a diagnosis. In order to simulate realistic constraints, the doctor agent is provided with a limited number of questions that they are able to ask the patient [14]. The doctor agent is also able to request test results from the measurement agent, specifying which test is to be performed (e.g. Chest X-Ray, EKG, blood pressure). When test results are requested, this also is counted toward the number of questions remaining.
Moderator agent
The function of the moderator is to determine whether the doctor agent has correctly diagnosed the patient at the end of the session. This agent is necessary because the diagnosis text produced by the doctor agent can be quite unstructured depending on the model, and must be parsed appropriately to determine whether the doctor agent arrived at the correct conclusion. For example, for a correct diagnosis of "Type 2 Diabetes Mellitus," the doctor might respond with the unstructured dialogue: "Given all the information we’ve gathered, including your symptoms, elevated blood sugar levels, presence of glucose and ketones in your urine, and unintentional weight loss I believe a diagnosis of Type 2 Diabetes with possible insulin resistance is appropriate," and the moderator must determine if this diagnosis was correct. This evaluation may also become more complicated, such as in the following example diagnosis: "Given your CT and blood results, I believe a diagnosis of PE is the most reasonable conclusion," where PE (Pulmonary Embolism) represents the correct diagnosis abbreviated.
@@
Figure 3: (Top) Demonstration of normalized accuracy ($\text{Accuracy}_{\text{bias}}$ / $\text{Accuracy}_{\text{No Bias}}$) in the presence of implicit and cognitive biases for both doctor and patient with GPT-4 (green) and Mixtral-8x7B (orange). GPT-4 accuracy was not susceptible to instructed biases, whereas Mixtral-8x7B was. Further results suggest GPT-4 rejects executing biases, whereas Mixtral-8x7B is more willing to represent bias (see section Bias and diagnostic accuracy). (Bottom) Ratings provided after diagnosis from GPT-4 patient agents with presented biases. While there were not large reductions in accuracy, biased patients had much less confidence in their treatment, lower compliance, and lower willingness for consultation. Left. Patient confidence in doctor. Middle. Patient compliance, indicating self-reported willingness to follow up with therapy. Right. Patient consultation rating, indicating willingness to consult with this doctor again.
@@
Previous work has indicated that LLMs can display racial biases[15] and might also lead to incorrect diagnoses due to inaccurate patient feedback[16]. Additionally, it has been found that the presence of prompts which induce cognitive biases can decrease the diagnostic accuracy of LLMs by as much as 26%[17]. The biases presented in this work intended to mimic cognitive biases that affect medical practitioners in clinical settings. However, these biases were quite simple, presenting a cognitive bias snippet at the beginning of each question (e.g. "Recently, there was a patient with similar symptoms that you diagnosed with permanent loss of smell"). This form of presentation did not allow for the bias to present in a realistic way, which is typically subtle and through interaction. We present biases that have been studied in other works from two categories: cognitive and implicit biases (Fig. 4). These are discussed below.
Cognitive biases
Cognitive biases are systematic patterns of deviation from norm or rationality in judgment, where individuals draw inferences about situations in an illogical fashion[18]. These biases can impact the perception of an individual in various contexts, including medical diagnosis, by influencing how information is interpreted and leading to potential errors or misjudgments. The effect that cognitive biases can have on medical practitioners is well characterized in literature on misdiagnosis[19]. In this work, we introduce cognitive bias prompts in the LLM system prompt for both the patient and doctor agents. For example, the patient agent can be biased toward believing their symptoms are pointing toward them having a particular disease (e.g. cancer) based on their personal internet research. The doctor can also be biased toward believing the patient symptoms are showing them having a particular disease based on a recently diagnosed patient with similar symptoms (recency bias).
Implicit biases
Implicit biases are associations held by individuals that operate unconsciously and can influence judgments and behaviors towards various social groups[20]. These biases may contribute to disparities in treatment based on characteristics such as race, ethnicity, gender identity, sexual orientation, age, disability, health status, and others, rather than objective evidence or individual merit. These biases can affect interpersonal interactions, leading to disparities in outcomes for the patient, and are well characterized in the medical literature[20, 21, 22]. Unlike cognitive biases, which often stem from inherent flaws in human reasoning and information processing, implicit biases are primarily shaped by societal norms, cultural influences, and personal experiences. In the context of medical diagnosis, implicit biases can influence a doctor’s perception, diagnostic investigation, and treatment plans for a patient. Implicit biases of patients can affect their trust—which is needed to open up during history taking—and their compliance with a doctor’s recommendations[21]. Thus, we define implicit biases for both the doctor and patient agents.
Our studied biases are shown in Figure 3. The bias prompt given to the agent is further discussed in the Appendix B.
@@
Figure 5: Accuracy of various doctor language models on AgentClinic-MedQA using GPT-4 patient and measurement agents (left). Accuracy of GPT-4 on AgentClinic-MedQA based on patient language model (middle). Accuracy on AgentClinic-MedQA by number of inferences (right).
@@
In order to build agents that are grounded in medically relevant situations, we use curated questions from the US Medical Licensing Exam (USMLE) and from the New England Journal of Medicine (NEJM) case challenges. These questions are concerned with diagnosing a patient based on a list of symptoms, which we use in order to build the Objective Structured Clinical Examination (OSCE) template that our agents are prompted with. For AgentClinic-MedQA, we first select from a random sample of 107 questions from the MedQA dataset and then populate a structured JSON formatted file containing information about the case study (e.g. test results, patient history) which is used as input to each of the agents. The exact structure of this file is demonstrated in Appendix C as well as an example case study shown in Appendix C. In general, we separate information by what is provided to each agent, including the objective for the doctor, patient history and symptoms for the patient, physical examination findings for the measurement, and the correct diagnosis for the moderator. We initially use an LLM (GPT-4) to populate the structured JSON, and then manually validate each of the case scenarios (Fig. 2). For AgentClinic-NEJM we select a curated sample of 15 questions from NEJM case challenges and proceed with the same template formatting as AgentClinic-MedQA.
Results
Comparison of models
@@ Results
Here we discuss the accuracy of various language models on AgentClinic-MedQA. We evaluate six models in total: GPT-4, GPT-4o, Mixtral-8x7B, GPT-3.5, Llama 3 70B-instruct, and Llama 2 70B-chat (discussed in detail in Appendix A). Each model acts as the doctor agent, attempting to diagnose the patient agent through dialogue. The doctor agent is allowed N=20 patient and measurement interactions before a diagnosis must be made. For this evaluation, we use GPT-4 as the patient agent for consistency. The accuracies of the models are presented in Figure 5: GPT-4 at 52%, GPT-4o and GPT-3.5 at 38%, Mixtral-8x7B at 37%, Llama 3 70B-instruct 30%, and Llama 2 at 70B-chat 9%.
@@
We also show results comparing the accuracy of these models on MedQA and AgentClinic-MedQA in Figure 6. Overall, while MedQA accuracy was only weakly predictive of accuracy on AgentClinic-MedQA. These results align with studies performed on medical residents, which show that the USMLE is poorly predictive of resident performance [23].
Bias and diagnostic accuracy

### v4

@@ Abstract
Evaluating large language models (LLM) in clinical scenarios is crucial to assessing their potential clinical utility. Existing benchmarks rely heavily on static question-answering, which does not accurately depict the complex, sequential nature of clinical decision-making. Here, we introduce AgentClinic, a multimodal agent benchmark for evaluating LLMs in simulated clinical environments that include patient interactions, multimodal data collection under incomplete information, and the usage of various tools, resulting in an in-depth evaluation across nine medical specialties and seven languages. We find that solving MedQA problems in the sequential decision-making format of AgentClinic is considerably more challenging, resulting in diagnostic accuracies that can drop to below a tenth of the original accuracy. Overall, we observe that agents sourced from Claude-3.5 outperform other LLM backbones in most settings. Nevertheless, we see stark differences in the LLMs’ ability to make use of tools, such as experiential learning, adaptive retrieval, and reflection cycles. Strikingly, Llama-3 shows up to 92% relative improvements with the notebook tool that allows for writing and editing notes that persist across cases. To further scrutinize our clinical simulations, we leverage real-world electronic health records, perform a clinical reader study, perturb agents with biases, and explore novel patient-centric metrics that this interactive environment firstly enables.
1 Introduction
@@
One of the primary goals in Artificial Intelligence (AI) is to build interactive systems that are able to solve a wide variety of problems. The field of medical AI inherits this aim, with the hope of making AI systems that are able to solve problems which can improve patient outcomes. Recently, many general-purpose large language models (LLMs) have demonstrated the ability to solve hard problems, some of which are considered challenging even for humans (Thirunavukarasu et al., 2023). Among these, LLMs have quickly surpassed the average human score on the United States Medical Licensing Exam (USMLE) in a short amount of time, from 38.1% in September 2021 (Gu et al., 2021) to 90.2% in November 2023 (Nori et al., 2023) (human passing score is 60%, human expert score is 87%  (Liévin et al., 2023)). While these LLMs are not designed to replace medical practitioners, they could be beneficial for improving healthcare accessibility and scale for the over 40% of the global population facing limited healthcare access (Organization et al., 2016) and an increasingly strained global healthcare system (McIntyre & Chow, 2020).
@@
However, there still remain limitations to these systems that prevent their application in real-world clinical environments. Recently, LLMs have shown the ability to encode clinical knowledge (Singhal et al., 2023; Vaid et al., 2023), retrieve relevant medical texts (Xiong et al., 2024), and perform accurate single-turn medical question-answering (Liévin et al., 2022; Nori et al., 2023; Wu et al., 2023; Chen et al., 2023). However, clinical work is a multiplexed task that involves sequential decision making, requiring the doctor to handle uncertainty with limited information and finite resources while compassionately taking care of patients and obtaining relevant information from them. This capability is not currently reflected in the static multiple choice evaluations (that dominate the recent literature) where all the necessary information is presented in a case vignettes and where the LLM is tasked to answer a question, or to just select the most plausible answer choice for a given question.
@@
In this work, we introduce AgentClinic, an open-source multimodal agent benchmark for simulating clinical environments. We improve upon prior work by simulating many parts of the clinical environment using language agents in addition to patient and doctor agents. Through the interaction with a measurement agent, doctor agents can perform simulated medical exams (e.g. temperature, blood pressure, EKG) and order medical image readings (e.g. MRI, X-ray) through dialogue. We also support the ability for agents to exhibit 24 different biases that are known to be present in clinical environments. We also present environments from 9 medical specialties, 7 different languages, and a study on incorporating various agent tools and reasoning techniques. Furthermore, our evaluation metrics go beyond diagnostic accuracy by giving emphasis to the patient agents with measures like patient compliance and consultation ratings.
Our key contributions are summarized as follows:
@@
    We challenge how large language and vision models should be evaluated for medical diagnosis with the introduction of AgentClinic. These diagnostic challenges are not static QAs, but are interactive, dialogue-driven, sequential decision making environments that require data collection, ordering appropriate medical exams, and understanding medical images across patients with unique family histories, lifestyle habits, age categories, and diseases.
@@
    A system for incorporating complex biases that can affect the dialogue and decisions of patient and doctor agents. We present results on diagnostic accuracy and patient perception for agents that are affected by cognitive and implicit biases. We find that doctor and patient biases can lower diagnostic accuracy, affect the patient’s willingness to follow through with treatment (compliance), reduce patient’s confidence in their doctor, and lower willingness for follow-up consultations.
@@
    We introduce patient agents built from real clinical cases sourced from electronic health record data, including an agent-based system for providing simulated medical exams (e.g. temperature, blood pressure, EKG) based on realistic disease test findings. We also introduce patient cases from nine medical specialties and seven multilingual environments to better support specialist applications and diverse language backgrounds. We also present realism and empathy ratings from clinicians for the resulting dialogue.
@@
    We allow doctor agents to use a variety of tools, such as browsing the web, textbooks, perform reflection cycles, take and edit notes in a notebook that persists over different patient scenarios. We show that current LLMs vastly differ in how much they benefit from these tools, with some models demonstrating large accuracy increases while others decrease in accuracy.
Figure 1: Running language agents in AgentClinic. (Left) Workflow diagram of agents in AgentClinic. The doctor agent interacts with tools and agents in order to arrive at a diagnosis. Moderator agent compares conclusion to ground truth diagnosis at the end of the simulation. (Right) Example dialogue between agents in the AgentClinic benchmark.
2 AgentClinic: a multimodal agent benchmark for clinical decision making
In this section we describe AgentClinic, which uses LLM agents to simulate a clinical environment.
@@ 2 AgentClinic: a multimodal agent benchmark for clinical decision making
Four language agents are used in the AgentClinic benchmark: a patient agent, doctor agent, measurement agent, and a moderator (Figure 1). Each language agent has specific instructions and is provided unique information that is only available to that particular agent. These instructions are provided to an LLM which carries out their particular role. The doctor agent serves as the model whose performance is being evaluated, and the other three agents serve to provide this evaluation. A detailed description of each agent is provided in Appendix A.1.
@@
Previous work has indicated that LLMs can display racial biases (Omiye et al., 2023) and might also lead to incorrect diagnoses due to inaccurate patient feedback (Ziaei & Schmidgall, 2023). Additionally, it has been found that the presence of prompts which induce cognitive biases can decrease the diagnostic accuracy of LLMs by as much as 26% (Schmidgall et al., 2024). The biases presented in this work intended to mimic cognitive biases that affect medical practitioners in clinical settings. However, these biases were quite simple, presenting a cognitive bias snippet at the beginning of each question (e.g. "Recently, there was a patient with similar symptoms that you diagnosed with permanent loss of smell"). This form of presentation did not allow for the bias to present in a realistic way, which is typically subtle and through interaction. We present clinically relevant biases that have been studied in other works from two categories: cognitive and implicit biases (Fig. 7). These biases are introduced by adding context into the agent’s system prompt instructing them to play out that bias as part of their role. For instance, to simulate sexual orientation bias, the patient agent receives the prompt: "You are uncomfortable with your doctor because you find out that they are a particular sexual orientation and you do not trust their judgement. This affects how you interact with them." These are discussed in detail in Appendix A.2.
@@
Figure 2: Accuracy of various doctor language models on AgentClinic-MedQA using GPT-4 patient and measurement agents (left). Accuracy of GPT-4 on AgentClinic-MedQA based on patient language model (middle). Accuracy on AgentClinic-MIMIC-IV by number of using GPT-4 patient and measurement agents (right).
@@
In order to build agents that are grounded in medically relevant situations, we use a random sample of diagnostic questions from the US Medical Licensing Exam (USMLE), from deidentified electronic health records (MIMIC-IV) (Johnson et al. (2023)), and from the New England Journal of Medicine (NEJM) case challenges. These questions are concerned with diagnosing a patient based on a list of symptoms, which we use in order to build the Objective Structured Clinical Examination (OSCE) template that our agents are prompted with. For AgentClinic-MedQA and AgentClinic-MIMIC-IV, we first select from a sample of questions from the MedQA and MIMIC-IV dataset respectively and then populate a structured JSON formatted file containing information about the case study (e.g. test results, patient history) which is used as input to each of the agents. The exact structure of this file is demonstrated in Appendix I as well as an example case study shown in Appendix J. In general, we separate information by what is provided to each agent, including the objective for the doctor, patient history and symptoms for the patient, physical examination findings for the measurement, and the correct diagnosis for the moderator. We initially use an LLM (GPT-4) to populate the structured JSON, and then manually validate each of the case scenarios. For AgentClinic-NEJM we select a curated sample of 120 questions from NEJM case challenges and proceed with the same template formatting as AgentClinic-MedQA/MIMIC-IV.
Multilingual and Specialist cases
Multilingual patient cases are converted from AgentClinic-MedQA to the the target language using GPT-4 and then manually corrected by native speakers. Agents are then prompted to perform dialogue in the target language. We chose to focus on six languages: Chinese, Hindi, Korean, Spanish, French, and Persian. The selection of these languages aims to address the need for medical AI systems capable of operating in multilingual healthcare environments. Specialist cases use case report questions from the MedMCQA dataset (Pal et al. (2022)). These questions include case reports from 20 different medical specialties, from which we chose to focus on 9 patient-focused specialties in AgentClinic-Spec: emergency medicine, geriatrics, pharmacology, internal medicine, psychiatry, ophthalmology, otolaryngology, and pediatrics.
3 Results
3.1 Comparison of models
@@
Here we discuss the accuracy of various language models on AgentClinic-MedQA. We evaluate seven models in total: Claude-3.5-Sonnet, GPT-4, GPT-4o, Mixtral-8x7B, GPT-3.5, Llama 3 70B-Instruct, and Llama 2 70B-chat (model details discussed in Appendix C). Each model acts as the doctor agent, attempting to diagnose the patient agent through dialogue. The doctor agent is allowed N=20 patient and measurement interactions before a diagnosis must be made. For this evaluation, we use GPT-4 as the patient agent for consistency. The accuracy of each models is presented in Figure 2: Claude-3.5 62.1% $\pm$ 3.3, GPT-4 at 51.6% $\pm$ 3.3, Mixtral-8x7B at 37.1% $\pm$ 3.1, GPT-3.5 at 36.6%, GPT-4o 34.2% $\pm$ 3.4 $\pm$ 3.2, Llama 3 70B at 19% $\pm$ 2.5, and Llama 2 at 70B-chat 4.5% $\pm$ 1.3. Confidence intervals for all experiments are provided in Appendix D.
Figure 3: Comparison of accuracy of models on MedQA and AgentClinic-MedQA. We find that MedQA accuracy is not predictive of accuracy on AgentClinic-MedQA.
We use the same configuration for AgentClinic-MIMIC-IV, with the accuracy of each model presented in Figure 2: Claude-3.5 achieves the highest accuracy at 42.9% $\pm$ 3.3, followed by GPT-4 at 34.0% $\pm$ 3.1. GPT-3.5 performed at 27.5% $\pm$ 3.0, while Mixtral-8x7B came in at 29.5% $\pm$ 3.1. GPT-4o demonstrated a lower accuracy of 24.0% $\pm$ 2.9, with Llama 3 70B scoring 8.5% $\pm$ 1.9, and Llama 2 70B-chat performing the worst at 13.5% $\pm$ 2.2.
We also find that the diagnostic accuracy in AgentClinic-MedQA is influenced by both the amount of interaction time and the choice of patient language model. Reducing the number of interactions from N=20 to N=10 significantly decreases accuracy from 52% to 25%, likely due to insufficient information being gathered, while increasing N beyond 20 to N=30 slightly reduces accuracy, possibly due to the complexity of processing larger inputs (Appendix F.1). Additionally, the choice of patient agent affects accuracy, with GPT-4 (52%) patient agents leading to higher diagnostic accuracy than GPT-3.5 (48%) or Mixtral (46%) agents, likely because GPT-4 provides more detailed responses (Appendix F.2). Interestingly, when a GPT-3.5 doctor interacts with a GPT-4 patient, accuracy is marginally higher than when both doctor and patient are GPT-3.5, which may suggest challenges in cross-model communication (Panickssery et al. (2024)).
@@
We also show results comparing the accuracy of these models on MedQA and AgentClinic-MedQA in Figure 3. Overall, MedQA accuracy was only weakly predictive of accuracy on AgentClinic-MedQA. These results align with studies performed on medical residents, which show that the USMLE is poorly predictive of resident performance (Lombardi et al., 2023).
3.2 How does bias affect the diagnostic accuracy of the doctor agent?
Figure 4: (Top) Demonstration of normalized accuracy ($\text{Accuracy}_{\text{bias}}$ / $\text{Accuracy}_{\text{No Bias}}$) with implicit and cognitive biases with GPT-4 (green) and Mixtral-8x7B (orange). GPT-4 accuracy was not susceptible to biases, whereas Mixtral-8x7B was. Results suggest GPT-4 rejects executing biases, whereas Mixtral-8x7B is more willing to represent bias (see section Bias and diagnostic accuracy). (Bottom) Ratings provided after diagnosis from GPT-4 patient agents with presented biases. While there were not large reductions in accuracy, biased patients had less confidence in their treatment, lower compliance, and lower willingness for consultation. Left. Patient confidence in doctor. Middle. Patient compliance, indicating self-reported willingness to follow up with therapy. Right. Patient consultation rating, indicating willingness to consult with this doctor again.